One card is drawn from a standard deck of cards. What is the probability that card is a jack, given that the card is red?
One card is drawn from a standard deck of cards. What is the probability that card is a heart, given that the card is red?

Answer by stanbon(75887)

One card is drawn from a standard deck of cards. What is the probability that card is a jack, given that the card is red?
"given that the card is red" means you should only consider red cards.
there are 2 red jacks, so
Ans:: 2/26 = 1/13
========================
One card is drawn from a standard deck of cards. What is the probability that card is a heart, given that the card is red?
# of red cards: 26
# of hearts that are red: 13
Ans: 13/26 = 1/2
======================
Cheers,
Stan H.
================